Chalcopyrite (CuFeS2) is a principal mineral of copper. Calculate the number of

kilograms of Cu in 3.71 X 103kg of chalcopyrite.


1
Expert's answer
2022-04-13T10:35:04-0400

Solution:

Calculate the mass percent of Cu in chalcopyrite:

%Cu = (Molar mass of Cu / Molar mass of CuFeS2) × 100%

The molar mass of Cu is 63.546 g/mol

The molar mass of CuFeS2 is 183.52 g/mol

Therefore,

%Cu = (63.546 / 183.52) × 100% = 34.63%


Calculate the mass of Cu in chalcopyrite:

Mass of CuFeS2 is 3.71×103 kg

Therefore,

Mass of Cu = %Cu × Mass of CuFeS2 / 100%

Mass of Cu = (34.63% × 3.71×103 kg) / 100% = 1284.77 kg

1284.77 kg of Cu = 1.28×103 kg of Cu


Answer: 1284.77 kg of Cu